# Latent Factor Analysis in Short Panels

### v2

@@ 2 Latent factor model and likelihood ratio test
This paper focuses mainly on testing hypotheses on the number of latent factors $k$ when $T$ is fixed and $n\rightarrow\infty$. The fixed $T$ perspective makes FA especially well-suited for applications with short panels. Indeed, we work conditionally on the realizations of the latent factors $F$ and treat their values as parameters to estimate. In comparison with the standard small $n$ and large $T$ framework in traditional asset pricing (e.g. Shanken (1992) with observable factors), here factors and loadings are interchanged in the sense that the $\beta_{i}$ and $F$ play the roles of the “factors" and the “factor loadings" in FA. We depart from classical FA since the $\beta_{i}$ are not considered as random effects with a Gaussian distribution but rather as fixed effects, namely incidental parameters. [^6] Chamberlain (1992) studies semiparametrically efficient estimation in panel models with fixed effects and short $T$ using moment restrictions from instrumental variables. Our approach does not rely on availability of valid instruments. Moreover, in Assumption 1, we neither assume Gaussianity nor we impose sphericity of the covariance matrix of the error terms. Besides we accommodate weak cross-sectional dependence and ARCH effects in idiosyncratic errors (see Section 3). Hence, the FA estimators defined below correspond to maximizers of a Gaussian pseudo likelihood. By-products of our analysis are the feasible asymptotic distributions of FA estimators of $F$ and $V_{\varepsilon}$ in more general settings than in the available literature (e.g.  Anderson and Amemiya (1988)), which we present in Appendix E.
The test statistic we consider for conducting inference on the number of latent factors $k$ is function of the elements of the symmetric matrix
$\hat{S}=\hat{V}_{\varepsilon}^{-1/2}M_{\hat{F},\hat{V}_{\varepsilon}}(\hat{V}_{y}-\hat{V}_{\varepsilon})M_{\hat{F},\hat{V}_{\varepsilon}}^{\prime}\hat{V}_{\varepsilon}^{-1/2},$ (2)
where $\hat{V}_{y}=\frac{1}{n}\tilde{Y}\tilde{Y}^{\prime}$ is the sample (cross-sectional) variance matrix (the $n$ columns of $\tilde{Y}$ are $y_{i}-\bar{y}$ and $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ is the vector of cross-sectional means), $M_{F,V}:=I_{T}-F(F^{\prime}V^{-1}F)^{-1}F^{\prime}V^{-1}$ is the Generalized Least Squares (GLS) projection matrix orthogonal to $F$ for variance $V$, and $\hat{F}$ and $\hat{V}_{\varepsilon}$ are the FA estimators computed under the assumption that there are $k$ latent factors. In the following, we use the same notation for the matrix-to-vector diag operator and the vector-to-matrix diag operator. Hence, $diag(A)$ for a matrix $A$ denotes the vector in which we stack the diagonal elements of matrix $A$, and $diag(a)$ for a vector $a$ denotes a diagonal matrix with the elements of $a$ on the diagonal. From Anderson (2003) Chapter 14, the FA estimators $\hat{F}$, $\hat{V}_{\varepsilon}$ maximize a Gaussian pseudo likelihood (Appendix E.1) and meet the first order conditions: [^7] The normalization in (FA2) applies for $\hat{\gamma}_{j}\geq 0$, which holds with probability approaching $1$. Otherwise, the first-order conditions of the FA estimators hold with $\hat{\gamma}_{j}$ replaced by its positive part; see Anderson (2003) for similar positivity constraints.
@@
$1+\hat{\gamma}_{j}$, $j=1,...,k$, normalized such that $\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=diag(\hat{\gamma}_{1},...,\hat{\gamma}_{k})$.
@@
The number of degrees of freedom is $df=\frac{1}{2}((T-k)^{2}-T-k)$. [^8] Integer $df$ is the number of different elements in data matrix $\hat{V}_{y}$, i.e., $\frac{1}{2}T(T+1)$, plus the number of normalization constraints $\frac{1}{2}k(k-1)$ in equations $F^{\prime}V_{\varepsilon}^{-1}F=diag$, minus the number of unknown parameters $(k+1)T$ (Anderson (2003)). In OA Table 3, we list the largest admissible number $k$ of latent factors as a function of $T$ such that $df\geq 0$. Generalizations to non-diagonal $V_{\varepsilon}$ are possible (Jöreskog (1970)), but stronger restrictions are required to maintain nonnegative degrees of freedom. Besides, numerical optimisation becomes less tractable. It is required that $df\geq 0$ for estimation, and we need $df>0$ to test the null hypothesis of $k$ latent factors (see Proposition 2 (a) below).
Statistic $\hat{S}$ in Equation (2) checks if the difference between the sample variance-covariance $\hat{V}_{y}$ and diagonal matrix $\hat{V}_{\varepsilon}$ is a symmetric matrix of reduced rank $k$, with range spanned by the range of $\hat{F}$. The probability limit of $\hat{S}$ is nil under the null hypothesis of $k$ latent factors. We get further insights from the next result.
Proposition 1
Under Assumption 1, (a) the eigenvalues of matrix $\hat{S}$ are: $\hat{\gamma}_{j}$, for $j=k+1,...,T$, and 00, with multiplicity $k$, where $1+\hat{\gamma}_{j}$ for $j=k+1,...,T$ are the $T-k$ smallest eigenvalues of $\hat{V}_{y}\hat{V}_{\varepsilon}^{-1}$, (b) the squared Frobenius norm is $\|\hat{S}\|^{2}=\sum_{j=k+1}^{T}\hat{\gamma}_{j}^{2}$, (c) $diag(\hat{S})=0$, and (d) we get
$\hat{S}=\hat{V}_{\varepsilon}^{-1/2}\left(\frac{1}{n}\hat{\varepsilon}\hat{\varepsilon}^{\prime}\right)\hat{V}_{\varepsilon}^{-1/2}-\hat{V}_{\varepsilon}^{-1/2}M_{\hat{F},\hat{V}_{\varepsilon}}\hat{V}_{\varepsilon}^{1/2},$ (3)
where $\hat{\varepsilon}=M_{\hat{F},\hat{V}_{\varepsilon}}\tilde{Y}$ is the $T\times n$ matrix of GLS residuals.
From Proposition 1 (d), we can interpret matrix $\hat{S}$ in terms of scaled cross-sectional averages of squared and cross-products of GLS residuals. In (3), we subtract $\hat{V}_{\varepsilon}^{-1/2}M_{\hat{F},\hat{V}_{\varepsilon}}\hat{V}_{\varepsilon}^{1/2}$ and not the identity because the residuals are orthogonal to $\hat{F}$ by construction. From Proposition 1 (c), the diagonal elements of matrix $\hat{S}$ vanish. Those elements are not informative for inference on the number of factors, and can be ignored when constructing the test statistics. This finding is natural because we expect that only the out-of-diagonal elements of $\frac{1}{n}\hat{\varepsilon}\hat{\varepsilon}^{\prime}$, i.e., the cross-sectional averages of cross-products of residuals for two different dates, are useful to check for omitted factors. [^9] The test in Connor and Korajczyk (1993) is built on cross-sectional averages of squared residuals, akin to diagonal terms of $\hat{S}$, but obtained by PCA instead of FA. However, their test statistic involves the difference of such cross-sectional averages for two consecutive dates, and relies on error sphericity.
@@
We now introduce the classical FA likelihood ratio (LR) statistic to test the null hypothesis $H_{0}(k)$ of $k$ latent factors:
${LR}(k):=-n\sum_{j=k+1}^{T}\log(1+\hat{\gamma}_{j}).$ (4)
The LR statistic in (4) only uses the information contained in the eigenvalues of matrix $\hat{S}$. Moreover, $LR(k)=\frac{n}{2}\|\hat{S}\|^{2}+o_{p}(1)$. [^10] Alternatively, we can use the squared norm statistic $\mathscr{T}(k)=n\|\hat{S}\|^{2}$. In the below, we focus on the LR test statistic since both deliver similar results in Monte Carlo experiments and in our empirics. It comes from a second-order expansion of the $\log$ function and the use of $\sum_{j=k+1}^{T}\hat{\gamma}_{j}=0$, from Proposition 1 (a) and (c) (see also Anderson (2003)), and $\sqrt{n}\hat{\gamma}_{j}=O_{p}(1)$, for $j=k+1,...,T$, from Propositions 1 and 2. Next, we establish feasibility of the asymptotic distribution of the LR statistic with $n\rightarrow\infty$ and $T$ fixed under a block-dependence structure.
3 Feasible asymptotic distributional theory
We start by defining the normalization for the latent factor matrix $F=[F_{1}:\cdots:F_{k}]$. Following classical FA, we set $\bar{\beta}=0$, $\tilde{V}_{\beta}=I_{k}$, and $F^{\prime}\tilde{V}_{\varepsilon}^{-1}F=diag(\gamma_{1},...,\gamma_{k})$, where $\bar{\beta}=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}$, $\tilde{V}_{\beta}=\frac{1}{n}\beta^{\prime}\beta$, and $\tilde{V}_{\varepsilon}=E[\frac{1}{n}\varepsilon\varepsilon^{\prime}]$. This normalization of the factor values is sample dependent, i.e., $F=F_{(n)}$. We skip index $n$ for the purpose of easing notation. Then, under our assumptions, we have $V_{y}:=\underset{n\rightarrow\infty}{\text{plim }}\hat{V}_{y}=FF^{\prime}+V_{\varepsilon}$, with $F^{\prime}V_{\varepsilon}^{-1}F=diag(\gamma_{1},...,\gamma_{k})$. [^11] The standardizations $\bar{\beta}=0$ and $\tilde{V}_{\beta}=I_{k}$ of the factor loadings wash out the incidental parameter problem (Neyman and Scott (1948); see Lancaster (2000) for a review) since the individual loadings do not appear in $V_{y}$. It explains why we are able to get consistent estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$ for large $n$ and fixed $T$. In particular, we have $V_{y}V_{\varepsilon}^{-1}F_{j}=(1+\gamma_{j})F_{j}$, i.e., the columns of $F$ are the eigenvectors of matrix $V_{y}V_{\varepsilon}^{-1}$ associated with eigenvalues $1+\gamma_{j}$, $j=1,...,k$ [^12] The remaining eigenvalue is equal to $1$ with multiplicity $T-k$. We have $F_{j}=\sqrt{\gamma_{j}}V_{\varepsilon}^{1/2}U_{j}$, where the $U_{j}$ are the orthonormal eigenvectors of $V_{\varepsilon}^{-1/2}V_{y}V_{\varepsilon}^{-1/2}$ for the $k$ largest eigenvalues $1+\gamma_{j}$.
We use a block-dependence structure to allow for weak cross-sectional dependence in errors.
@@ 3 Feasible asymptotic distributional theory
The block-dependence structure as in Assumption 2 is satisfied, for instance, when there are unobserved industry-specific factors independent among industries and over time, as in Ang, Liu, and Schwarz (2020). In empirical applications, blocks in $\Sigma$ can match industrial sectors (Fan, Furger, and Xiu (2016), Gagliardini, Ossola, and Scaillet (2016)). As already remarked, the diagonal elements of $V_{\varepsilon}$ are the sample realizations of the common component driving the variance of the error terms at times $t=1,...,T$; see e.g. Barigozzi and Hallin (2016), Renault, Van Der Heijden and Werker (2023) for theory and empirical evidence pointing to variance factors. A sphericity assumption cannot accommodate such a common time-varying component. Assumption 2 (a) is coherent with Assumption 1. Indeed, $\tilde{V}_{\varepsilon}=V_{\varepsilon}^{1/2}\frac{1}{n}\sum_{i,j=1}^{n}\sigma_{i,j}E[w_{i}w_{j}^{\prime}]V_{\varepsilon}^{1/2}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}V_{\varepsilon}$ is diagonal. Hence, $\tilde{V}_{\varepsilon}$ is a scalar multiple of $V_{\varepsilon}$, and converges to $V_{\varepsilon}$ under the normalization $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}=1$. That normalization is without loss of generality by rescaling of the parameters. Assumption 2 (c) builds on Bickel and Levina (2008), and $\delta<1$ holds under sparsity, vanishing correlations or mixing dependence within blocks. With blocks of equal size, Assumption 2 (d) holds for $J_{n}=n^{\bar{\alpha}}$ and $\bar{\alpha}>\frac{2\delta}{2\delta+1}$. Having $\delta<1$ helps relaxing this condition on block granularity, however it is not strictly necessary because we allow value $\delta=1$.

### v3

@@ 2 Latent factor model and likelihood ratio test
This paper focuses mainly on testing hypotheses on the number of latent factors $k$ when $T$ is fixed and $n\rightarrow\infty$. The fixed $T$ perspective makes FA especially well-suited for applications with short panels. Indeed, we work conditionally on the $T$ realizations of the latent factors $F$ and treat their values as $T\times k$ parameters to estimate. In comparison with the standard small $n$ and large $T$ framework in traditional asset pricing (e.g. Shanken (1992) with observable factors), here factors and loadings are interchanged in the sense that the $\beta_{i}$ and $F$ play the roles of the “factors" and the “factor loadings" in FA. We depart from classical FA since the $\beta_{i}$ are not considered as random effects with a Gaussian distribution but rather as fixed effects, namely non-random incidental parameters. Moreover, in Assumption 1, we neither assume Gaussianity nor do we impose sphericity of the covariance matrix of the error terms. In addition, we accommodate weak cross-sectional dependence and general forms of heteroskedasticity (individual and common across assets) in idiosyncratic errors (see Section 3). Hence, the FA estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$ defined below correspond to maximizers of a Gaussian pseudo likelihood. By-products of our analysis are the feasible asymptotic distributions of FA estimators of $F$ and $V_{\varepsilon}$ in more general settings than in the available literature (e.g. Anderson and Amemiya (1988)), which we present in Appendix B. Chamberlain (1992) studies semiparametrically efficient estimation in panel models with fixed effects and short $T$ using moment restrictions from instrumental variables. Our approach does not rely on availability of valid instruments. Another approach to deal with fixed $T$ is to increase the sampling frequency. Andersen et al. (2025) develop a test for determining the number of jump factors under fixed $T$, large $n$, and increasing sampling frequency asymptotics. In their framework, the idiosyncratic noise vanishes asymptotically and therefore does not need to be estimated.
Let $\hat{V}_{y}=\frac{1}{n}\tilde{Y}\tilde{Y}^{\prime}$ be the sample $T\times T$ cross-sectional covariance matrix, where the $n$ columns of $\tilde{Y}$ are $y_{i}-\bar{y}$ and $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ is the vector of cross-sectional means. The covariance $\hat{V}_{y}$ is a sufficient statistic for FA estimation of the parameter $\theta=(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime})^{\prime}$. In the following, we use the same notation for the matrix-to-vector diag operator and the vector-to-matrix diag operator. Hence, $diag(A)$ for a matrix $A$ denotes the vector in which we stack the diagonal elements of matrix $A$, and $diag(a)$ for a vector $a$ denotes a diagonal matrix with the elements of $a$ on the diagonal. Moreover, to normalize the latent factor matrix $F=[F_{1}:\cdots:F_{k}]$, we follow classical FA and set $\bar{\beta}=0$, $\tilde{V}_{\beta}=I_{k}$, and $F^{\prime}\tilde{V}_{\varepsilon}^{-1}F=diag(\gamma_{1},...,\gamma_{k})$, where $\bar{\beta}=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}$ and $\tilde{V}_{\beta}=\frac{1}{n}\beta^{\prime}\beta$ are the sample mean and second moment of the betas (we treat $\beta_{i}$ as fixed effects, and not random variables), and $\tilde{V}_{\varepsilon}=E[\frac{1}{n}\varepsilon\varepsilon^{\prime}]$ is the finite-$n$ analog of $V_{\varepsilon}$. This normalization of the factor values is sample dependent, i.e., $F=F_{(n)}$. We suppress index $n$ for the purpose of easing notation. The normalization can be made without loss of generality and serves to remove some of the fundamental indeterminacy that arises when working with latent factors (see Anderson, 2003, Chapter 14). Then, under our assumptions, we have $V_{y}:=\underset{n\rightarrow\infty}{\text{plim }}\hat{V}_{y}=FF^{\prime}+V_{\varepsilon}$, with $F^{\prime}V_{\varepsilon}^{-1}F=diag(\gamma_{1},...,\gamma_{k})$. In particular, we have $V_{y}V_{\varepsilon}^{-1}F_{j}=(1+\gamma_{j})F_{j}$, i.e., the columns of $F$ are the eigenvectors of matrix $V_{y}V_{\varepsilon}^{-1}$ associated with eigenvalues $1+\gamma_{j}$, j=1,,kj=1,...,k. The standardizations $\bar{\beta}=0$ and $\tilde{V}_{\beta}=I_{k}$ wash out the incidental parameter problem (Neyman and Scott (1948); see Lancaster (2000) for a review) since the individual loadings do not appear in $\mu_{y}:=\underset{n\rightarrow\infty}{\text{plim}}~\bar{y}$ nor in $V_{y}$ and we do not need to estimate them. It explains why we are able to get consistent estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$ for large $n$ and fixed $T$.
As in Anderson (2003) Chapter 14 (Section 14.3.1), the FA estimators $\hat{F}$, $\hat{V}_{\varepsilon}$ with $k$ latent factors maximize a Gaussian pseudo likelihood written under Assumption 1 (Appendix B.1). Even if we write the Gaussian PML under the assumption that $\beta_{i}$ are i.i.d. $N(0,I_{k})$ in Appendix B.1, we do not assume that to derive our asymptotic results. When both $n$ and $T$ are large and cross-sectional independence holds, Bai and Li (2012) show that the factor estimates given by a Gaussian PML approach and a Generalized Least Squares (GLS) approach (weighted PCA), albeit different, share the same asymptotic distribution. Our factor estimates driven by the first order conditions (FA1) and (FA2) do not share the same asymptotic distribution as the ones given by a GLS approach since the GLS estimates build on factor loadings estimates and thus suffer from a bias due to the incidental parameter problem under large $n$ and fixed $T$.
The FA estimators meet the first order conditions:
@@
$1+\hat{\gamma}_{j}$, j=1,,kj=1,...,k, normalized such that $\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F}=diag(\hat{\gamma}_{1},...,\hat{\gamma}_{k})$.
@@
The normalization in (FA2) applies for $\hat{\gamma}_{j}\geq 0$, which holds with probability approaching $1$. Otherwise, the first-order conditions of the FA estimators hold with $\hat{\gamma}_{j}$ replaced by its positive part. If error variance $V_{\varepsilon}$ where known, i.e., $\hat{V}_{\varepsilon}=V_{\varepsilon}$, condition (FA2) yields that $V_{\varepsilon}^{-1/2}\hat{F}$ is the matrix of the orthogonal eigenvectors of $V_{\varepsilon}^{-1/2}\hat{V}_{y}V_{\varepsilon}^{-1/2}$. It is akin to PCA applied on pre-whitened data $V_{\varepsilon}^{-1/2}Y$, up to normalization. Because $V_{\varepsilon}$ is unknown and needs to be estimated, condition (FA1) is an implicit equation that yields $\hat{V}_{\varepsilon}$ from the diagonal of $\hat{V}_{y}-\hat{F}\hat{F}^{\prime}$. The numerical calculation of the FA estimates from (FA1) and (FA2) is described in Magnus and Neudecker (2007) Chapter 17. The number of degrees of freedom is $df=\frac{1}{2}((T-k)^{2}-T-k)$. It corresponds to the number of different elements in data matrix $\hat{V}_{y}$, i.e., $\frac{1}{2}T(T+1)$, plus the number of normalization constraints $\frac{1}{2}k(k-1)$ in equations $F^{\prime}V_{\varepsilon}^{-1}F=diag$, minus the number of unknown parameters $(k+1)T$ (Anderson (2003)). It is required that $df\geq 0$ for estimation, and we need $df>0$ to test the null hypothesis of $k$ latent factors (see Proposition 1 (a) below). In OA Table 3, we list the largest admissible number $k$ of latent factors as a function of $T$ such that $df\geq 0$.
An alternative to maximum likelihood estimation is minimum distance estimation, also called Asymptotically Distribution-Free (ADF) estimation in covariance structure analysis (see Zheng and Bentler (2024) for a comparison of the two approaches). The two approaches are related, since we can show that the second-order expansion of the Gaussian pseudo likelihood criterion underlying the FA estimator yields (minus) a minimum distance criterion with weighting matrix $(V_{y})^{-1}\otimes(V_{y})^{-1}$. The asymptotic properties of minimum distance estimation under block-dependence are not available in the literature and require additional developments.
@@
We conclude this section by introducing the classical FA likelihood ratio (LR) statistic to test the null hypothesis $H_{0}(k)$ of $k$ latent factors against the alternative hypothesis $H_{1}(k)$ of more than $k$ latent factors :
${LR}(k):=-n\sum_{j=k+1}^{T}\log(1+\hat{\gamma}_{j}),$ (2)
where the $\hat{\gamma}_{j}$ for j=k+1,,Tj=k+1,...,T are the TkT-k smallest eigenvalues of $\hat{V}_{y}\hat{V}_{\varepsilon}^{-1}-I_{T}$. We have $LR(k)=\frac{n}{2}\sum_{j=k+1}^{T}\hat{\gamma}_{j}^{2}+o_{p}(1)$, which results from a second-order expansion of the $\log$ function and the use of $\sum_{j=k+1}^{T}\hat{\gamma}_{j}=0$, from Lemma 1 in Appendix A and Anderson (2003), and $\sqrt{n}\hat{\gamma}_{j}=O_{p}(1)$, for j=k+1,,Tj=k+1,...,T, from the proof of Proposition 1 below. Hence, the LR statistic checks if the TkT-k smallest eigenvalues of $\hat{V}_{y}\hat{V}_{\varepsilon}^{-1}$ are close to $1$, which mimics the properties of the population quantities. Prewhitening $Y$ with $V_{\varepsilon}^{-1/2}$, gives limit variance $V_{\varepsilon}^{-1/2}V_{y}V_{\varepsilon}^{-1/2}=V_{\varepsilon}^{-1/2}FF^{\prime}V_{\varepsilon}^{-1/2}+I_{T}$. Hence, the $k$ largest values of $V_{y}V_{\varepsilon}^{-1}$ are above $1$ while the subsequent ones are equal to 1. It is the scree plot (plot of eigenvalues in descending order introduced by Cattell (1966)) related to $V_{y}V_{\varepsilon}^{-1}$ that is informative on the number of factors and not the one related to $V_{y}$ as in standard PCA. When sphericity fails, the latter should not be used in practice with fixed $T$ since the shape of the spectrum of $V_{y}V_{\varepsilon}^{-1}$ is not similar to the shape of the spectrum of $V_{y}$. Since the check is on the TkT-k smallest eigenvalues being close to 1, we cannot interpret the LR test as a test of the null hypothesis that the number of factors is less than or equal to $k$ and is thus different from the proposal by Chen and Fang (2019) for inference on the rank of a matrix. In Section 4, we establish feasibility of the asymptotic distribution of the LR statistic with $n\rightarrow\infty$ and $T$ fixed under a block-dependence structure.
3 Regularity assumptions
In this section, we list and comment the additional assumptions used to derive the large sample properties of the estimators and test statistics. We often denote by $C>0$ a generic constant, and we use $\delta_{j}(A)$ to denote the $j$th largest eigenvalue of a symmetric matrix $A$. Set $\Theta$ is a compact subset of $\{\theta=(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime})^{\prime}\in\mathbb{R}^{r}~:~V_{\varepsilon}~\text{is diagonal and positive definite},~F^{\prime}V_{\varepsilon}^{-1}F~\text{is diagonal, with}$$\text{diagonal elements ranked in decreasing order}\}$ with $r=(T+1)k$, and function $L_{0}(\theta)=\penalty-10000-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(V_{y}\Sigma(\theta)^{-1}\right)$ is the population FA criterion, where $\Sigma(\theta)=FF^{\prime}+V_{\varepsilon}$ and $V_{y}=\underset{n\rightarrow\infty}{\text{plim}}~\hat{V}_{y}$. Further, $\theta_{0}=(vec(F_{0})^{\prime},diag(V_{\varepsilon}^{0})^{\prime})^{\prime}$ denotes the vector of true parameter values under $H_{0}(k)$ and is an interior point of set $\Theta$. For a $T\times T$ symmetric matrix $A=(a_{i,j})$, let us define the $\frac{1}{2}T(T+1)\times 1$ vector $vech(A)=\left(\frac{1}{\sqrt{2}}a_{11},...,\frac{1}{\sqrt{2}}a_{T,T},\{a_{i,j}\}_{i<j}\right)^{\prime}$, where the pairs of indices $(i,j)$ with $i<j$ are ranked as (1,2),(1,3),,(1,T),(2,3),,(T1,T)(1,2),(1,3),...,(1,T),(2,3),...,(T-1,T). This definition of the half-vectorization operator for symmetric matrices differs from the usual one by the ordering of the elements, and the rescaling of the diagonal elements. It is more convenient for our purposes (see proof of Lemma 11). For instance, it holds $\frac{1}{2}\|A\|^{2}=vech(A)^{\prime}vech(A)$. Below, $\odot$ denotes the Hadamard product (i.e., element-wise matrix product), and $A^{\odot 2}:=A\odot A$.
@@ 3 Regularity assumptions
Assumption 3
Assumption 4
The loadings are normalized such that $\bar{\beta}=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}=0$ and $\tilde{V}_{\beta}:=\frac{1}{n}\sum_{i=1}^{n}\beta_{i}\beta_{i}^{\prime}\penalty-10000=I_{k}$, for any $n$. Moreover, $|\beta_{i}|\leq C$, for all $i$.
Assumption 5
We have $E[w_{i,t}^{8}]\leq C$ and $|\sigma_{i,j}|\leq C$, for all $i,j,t$.
Assumption 6
Under the null hypothesis $H_{0}(k)$, we have: $\Sigma(\theta)=\Sigma(\theta_{0})$, $\theta\in\Theta$ $\Rightarrow$ $\theta=\theta_{0}$, up to sign changes in the columns of $F$.
Assumption 7
Matrix $M_{F_{0},{V}^{0}_{\varepsilon}}^{\odot 2}$ is non-singular.
Assumption 8
Assumption 9
Assumption 10
@@
We use the block-dependence structure in Assumption 2 to allow for weak cross-sectional dependence in errors and get implementable estimators of population (unknown) quantities, i.e., feasible asymptotic results under fixed $T$, as stated in Proposition 1 and discussed in Section 4 below. Under large $n$ and $T$, other estimation strategies for the asymptotic variance are available (see e.g. Bai and Ng (2006a), Maldonado and Ruiz (2021), Kim (2022)). The modeling of the errors in Assumption 2 (a) is similar to Onatski (2010) and Ahn and Horenstein (2013) but with more restrictive structure because of the diagonality of $V_{\varepsilon}$ and the blocks in $\Sigma$. The block-dependence structure as in Assumption 2 (b) is satisfied, for instance, when there are unobserved industry-specific factors independent among industries and over time, as in Ang, Liu, and Schwarz (2020). In empirical applications, blocks in $\Sigma$ can match industrial sectors (Fan, Furger, and Xiu (2016), Gagliardini, Ossola, and Scaillet (2016)). As already remarked, the diagonal elements of $V_{\varepsilon}$ are the sample realizations of the common component driving the variance of the error terms at times t=1,,Tt=1,...,T; see e.g. Barigozzi and Hallin (2016), Renault, Van Der Heijden and Werker (2023) for theory and empirical evidence pointing to variance factors. A sphericity assumption cannot accommodate such a common time-varying component. Assumption 2 (a) is coherent with Assumption 1. Indeed, $\tilde{V}_{\varepsilon}=V_{\varepsilon}^{1/2}\frac{1}{n}\sum_{i,j=1}^{n}\sigma_{i,j}E[w_{i}w_{j}^{\prime}]V_{\varepsilon}^{1/2}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}V_{\varepsilon}$ is diagonal. Hence, $\tilde{V}_{\varepsilon}$ is a scalar multiple of $V_{\varepsilon}$, and converges to $V_{\varepsilon}$ under the normalization $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}=1$. That normalization is without loss of generality by rescaling of the parameters. Assumption 2 (c) builds on Bickel and Levina (2008), and $\delta<1$ holds under sparsity, vanishing correlations or mixing dependence within blocks. Assumption 2 (d) implies a minimum degree of granularity on the blocks, i.e., largest blocks cannot be too large and the number of blocks should grow fast enough. With blocks of equal size, Assumption 2 (d) holds for $J_{n}=n^{\bar{\alpha}}$ and $\bar{\alpha}>\frac{2\delta}{2\delta+1}$. Having $\delta<1$ helps relaxing this condition on block granularity, however it is not strictly necessary because we allow value $\delta=1$.
Assumption 3 removes the rotational indeterminacy (up to sign) that occurs when identifying the columns of matrix $F$ as eigenvectors of $V_{y}V^{-1}_{\varepsilon}$. Assumptions 4 and 5 require uniform bounds on factor loadings as well as on covariances and higher-order moments of the idiosyncratic errors. We can weaken Assumption 5 to require only existence of error moments slightly above order 4 (finite kurtosis) at the expense of further technicalities and more granular blocks in Assumption 2. Assumption 6 implies global identification in the FA model (see Lemma 5). Assumptions 4-6 yield consistency of FA estimators (see proof of Lemma 6). Assumption 7 is the local identification condition in the FA model (see Lemma 7). We use Assumption 8 together with Assumption 5 to invoke a CLT based on a multivariate Lyapunov condition (see proof of Lemma 2) to establish the asymptotic distribution of the LR statistic. To ease the verification of the Lyapunov condition, we bound a fourth-order moment of squared errors, which explains why we require finite eight-order moments in Assumption 5. We could relax this condition at the expense of a more sophisticated proof of Lemma 2. The mild Assumption 8 (b) requires that the smallest eigenvalue of $V[vech(w_{i}w_{i}^{\prime})]$ is bounded away from 0 for all assets $i$ up to a small fraction. In Assumption 8 (c), in order to have $\kappa_{n}$ bounded, we need either mixing dependence in idiosyncratic errors within blocks, i.e., $|\sigma_{i,j}|\leq C\rho^{|i-j|}$ for $i,j\in I_{m}$ and $0\leq\rho<1$, or vanishing correlations, i.e., $|\sigma_{i,j}|\leq Cb_{m,n}^{-\bar{s}}$ for all $i\neq j\in I_{m}$ and a constant $\bar{s}\geq 1/2$, with blocks of equal size. In Assumption 9, part (a) ensures a well-defined pseudo-true parameter value (White (1982)) under the alternative hypothesis. Since we maximize the likelihood over a compact set, a maximizer exists. If the maximizer $\theta^{*}$ is unique w.r.t. $V^{*}_{\varepsilon}$, and the first $k$ eigenvalues of $V_{y}(V^{*}_{\varepsilon})^{-1}-I_{T}$ are distinct (as in Assumption 3), then the pseudo-true values of the factors $F^{*}$ are unique up to column sign by (FA2). Part (b) is used to establish the consistency of the LR test under global alternative hypotheses (see proof of Proposition 1). Finally, Assumption 10 is used to apply a Lyapunov CLT (see proof of Lemma 8) when deriving the asymptotic normality of the FA estimators.
4 Feasible asymptotic distributional theory